OpenStudy (anonymous):

A model of a 50ft building was made using a scale of 2in:10 ft . What is the height of the model ????

OpenStudy (anonymous):

You'd set a proportion: 2 in/10 ft=x in/50 ft Cross multiply: 2*50=10*x 100=10x 10in=x